elizabeth is 68 years old and her son is 46 years old how many years ago was he 1/3 as old as she
Found 3 solutions by stanbon, Alan3354, ikleyn:
Answer by stanbon(75887)

elizabeth is 68 years old and her son is 46 years old how many years ago was he 1/3 as old as she
------
Let the answer be "x".
-----
Equation:
46-x = (1/3)(68-x)
138-3x = 68-x
-----
2x = 70
---
x = 35 yrs
--------------
Cheers,
Stan H.

Answer by Alan3354(69443)

The difference is always 22 years.
--> 33 & 11

Answer by ikleyn(52906)

.
46 - x = 3*(68-x).


Solve for x.

*******
    OR
*******

The difference is 68 - 46 = 22.


Hence, "then" he was 11, she was 33.


46-11 = 35 years ago.


Check.  68 - 33 = 35.   ! Correct !

Solved.